A ball is kicked horizontally at 8.0 m/s from a cliff 80m high. How far from the base of the cliff will the stone strike the ground?


1
Expert's answer
2020-10-27T11:17:00-0400

Let's first write the equations of motion of the ball in horizontal and vertical directions:


"x=v_{0x}t,""y=v_{0y}t+\dfrac{1}{2}gt^2,"

here, "x" is the horizontal displacement of the ball (or the distance from the base of the cliff to the ball), "v_{0x}=8.0\ \dfrac{m}{s}" is the initial horizontal velocity of the ball, "t" is the time that the ball takes to reach the ground, "y=80\ m" is the vertical displacement of the ball (or the height of the cliff), "v_{0y}=0\ \dfrac{m}{s}" is the initial vertical velocity of the ball and "g=9.8\ \dfrac{m}{s^2}" is the acceleration due to gravity.

Let's find the time that the ball takes to reach the ground from the second equation:


"t=\sqrt{\dfrac{2y}{g}}."

Finally, we can substitute "t" into the first equation and find the distance from the base of the cliff to the ball:


"x=v_{0x}t=v_{0x}\sqrt{\dfrac{2y}{g}},""x=8.0\ \dfrac{m}{s}\cdot\sqrt{\dfrac{2\cdot 80\ m}{9.8\ \dfrac{m}{s^2}}}=32\ m."

Answer:

"x=32\ m."